40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 20 14:42:34 2011, steve, Update, SUS, oplev gains today 
    Reply  Wed Jul 20 19:24:12 2011, rana, Update, SUS, oplev gains today Untitled.pnga.png
       Reply  Wed Jul 20 19:48:03 2011, Jamie, Update, SUS, Re: oplev gains today 
       Reply  Thu Jul 21 09:04:59 2011, valera, steve, Update, SUS, oplev gains were not optimized 
       Reply  Mon Jul 25 00:35:44 2011, rana, Update, SUS, something wrong with ETMY LR sensor a.pngbad.png
          Reply  Tue Jul 26 11:38:45 2011, steve, Update, SUS, ETMY bias checked 
    Reply  Thu Jul 21 16:06:35 2011, steve, Update, SUS, oplev gains today checked by conlog 
Message ID: 5004     Entry time: Wed Jul 20 19:24:12 2011     In reply to: 5001     Reply to this: 5005   5010   5025
Author: rana 
Type: Update 
Category: SUS 
Subject: oplev gains today 

I guess Valera forgot to elog it. Steve, please email him and get the info.

I started to check out the OL servos today so that our whole interferometer is not too floppy.

  • The ETMX OPLEV DAQ channels were not in the list. Jamie ran the activateDQ.py script and it came back. Right now, we have no diagnostics to know if people have run this or not so the frames will have missing data now and again depending on how forgetful the rebooters are. Perhaps we can get activateDQ put into the make file???
  • I turned ON all of the offset buttons on the OL1, etc. filter banks. This allows for the dark offsets to be set for the OL quadrants. With these buttons off it doesn't make any sense.
  • I noticed that there are white (INVALID) fields all over the OPTLEV_SERVO screens. This is just because the new SUS models have not captured all of the functionality of the old system. Needs fixing.

Untitled.png

Some of these OL spectra are not like the others...

 a.png

ELOG V3.1.3-